Calculate the number of normal modes of vibration of brf5 and xef2
1
Expert's answer
2016-09-08T10:55:04-0400
BrF5
[3n1-5] + [3n2-6]
3n-5 is the formula for number of modes of vibration in linear molecule
3n-6 is the formula for number of modes of vibration in non-linear
combined both because the given molecule is complex (BrF5)
where n= no. of electrons
where n1= no of electrons in linear (i.e., F & Lonepair in axial)
where n2= no of electrons in non-linear {4 F's in sq.plane}
F has 6e's
=> 3·(2+6)-5 + 3·(4·6)-6 = 3·8+3·24-11 = 24+72-11= 85
Modes of Vibration in BrF5= 85

XeF2 {also a complex} trigonal bipyramidal shape
lone pairs occupying a triangular plane
F's in axial because they're more EN than Lone Pair
=> 3·(2·6)-5 + 3·(3·2)-6 = 3·12+3·6-11= 36+18-11= 63
Modes of Vibration in XeF2=63